Social Science: In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate recieive?
Thanks.

Answer by ptaylor(1661)


Let x=number of votes loser got
Then x+220=number of votes the winner got
Now we are told that:
x+x+220=810 subtract 220 from both sides
2x=810-220
2x=590
x=295-----------------------------number of votes for loser
x+220=295+220=515------------------number of votes for winner

Ck
295+515=810
810=810

Hope this helps-----ptaylor